OpenStudy (anonymous):

A kilobyte is 210. A megabyte is 220 bytes. How many kilobytes are in a megabyte? Use the exponent rule for division to find your answer and give your answer in exponential form.

OpenStudy (jamesj):

What you have written makes no sense. A kilobyte is 1000 or sometimes 1024 bytes. And a megabyte is 1,000,000 bytes (or sometimes 1024^2 bytes).

OpenStudy (anonymous):

2^10 nto 210 and 2^20 not 220 ..

OpenStudy (jamesj):

Right ... now it makes sense. So 1 MB / 1 KB = 2^20/2^10 = 2^20 . 2^(-10) = 2^(20 - 10) = 2^10 Hence the number of KB per MB is ....

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

MB = megabyte, KB = kilobyte